The ethyl alcohol content of many beers produced in the Philippines is 4.05% (vol/vol). If the density of ethyl alcohol (FM=46.058) at room temperature is 0.7893 g/mL, what is the percent of ethyl alcohol in beer expressed as percent (wt/vol)?

A. 3.20% ethyl alcohol (wt/vol)

B. 5.13% ethyl alcohol (wt/vol)

C. 7.80% ethyl alcohol (wt/vol)

D. none of the above
Answer #1

4.05% v/v means , 100 mL of solution contains 4.05 mL ethyl alcohol.

Mass of solution = v×d

mass(m) = 4.05×0.7893 => 3.20

%wt/v = 3.20×100/3.20

= 3.20 % wt./vol